OpenStudy (unklerhaukus):

The Product Rule Which is more correct? \[\left(fg\right)^\prime=f^\prime g+g^\prime f\] or \[\left(fg\right)^\prime=f^\prime g+f g^\prime \]

OpenStudy (anonymous):

second one?

OpenStudy (unklerhaukus):

justify,

OpenStudy (blockcolder):

How can one be 'more' correct than the other one if they're the same?

OpenStudy (unklerhaukus):

they are not the same the ordering is different

OpenStudy (blockcolder):

The ordering is different but by commutative law of multiplication, they're the same.

OpenStudy (unklerhaukus):

can you assume commutativity ?

OpenStudy (unklerhaukus):

perhaps they are anti-commutative

OpenStudy (blockcolder):

Well, commutative law of multiplication states that \[\forall a,b\in \mathbb{R}\ (ab=ba)\] A function f from R to R (subsets of R will work too) goes like this: \[\begin{align} f:\mathbb{R}\ &\to \mathbb{R}\\ x\ &\mapsto f(x) \end{align}\] Thus, the assignment assigns a real number to another real number, which satisfies the commutative law of multiplication.

OpenStudy (anonymous):

both will give same result!!

OpenStudy (unklerhaukus):

not all products are commutative

OpenStudy (anonymous):

example??

OpenStudy (blockcolder):

Yeah, like the cross product.

OpenStudy (unklerhaukus):

matrices

OpenStudy (anonymous):

this is not cross product and this is not matrix!!

OpenStudy (unklerhaukus):

\[f,g \cancel \in \mathbb R\]

OpenStudy (blockcolder):

Personally, I'd like to memorize the rule in the 2nd form, because it helps me facilitate the memorization of this rule: \[\frac{d}{dt}[\mathbf{u}(t)\times\mathbf{v}(t)]=\mathbf{u}'(t)\times\mathbf{v}(t)+\mathbf{u}(t)\times\mathbf{v}'(t)\]

OpenStudy (blockcolder):

So how do the assignments f and g go, if they're not from R to R, such that the elements of the codomain won't satisfy commutativity?

OpenStudy (anonymous):

doing it any vch way vl not distinct the answer!!

OpenStudy (anonymous):

lol briefsssss

OpenStudy (anonymous):

I think the most easy is this one: \[\large (fg)' = fg' + gf'\]